A 1.5 MVA ,3.3 kV, 40 pole, three-phase star connected asynchronous alternator has reactance of Xd = 4.01Ω/phase and Xq = 2.88Ω/phase.For a lagging power factor of 0.8 when delivering full load. Calculate a developed power: 


1
Expert's answer
2021-08-28T06:16:10-0400

"V_t= \frac{3300}{\sqrt3}=1905.25 \\
\delta = \phi -\theta = 14.904\\
E_f= V_t \cos \delta +I_dX_d= 1905.25 \cos(14.9) +215.925*4.0=2707.0472 V\\"

"Developed \space power = \frac{E_f V_t}{X_d} \sin \delta + \frac{V_t^2}{2}(\frac{1}{X_2}-\frac{1}{X_d})\sin2 \delta\\
= \frac{4.688*3.3}{4.01} \sin 14.9 + \frac{3.3^2}{2}(\frac{1}{2.88}-\frac{1}{4.01})\sin2*14.9\\
=1.256 MW"